OpenStudy (anonymous):

3. Solve the quadratic equation y2 - 5y = -4 by factoring. A. {-2, 2} B. {2, 2} C. {4, 1} also, can you explain in as much detail as you can how that is the answer?

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

@charlotteakina which class are u in

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

y2 -5y = -4 y2 - 5y + 4 = 0 y2 - 4y - y + 4 = 0 y (y - 4) -1(y - 4) (y - 4)(y - 1) y = 4 , 1

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

so bye @charlotteakina today is my english exam so need to sleep I am from india :)